Status Report

NOAA GOES SATOPS Morning Report 8 Jan 2001

By SpaceRef Editor
January 8, 2001
Filed under ,

SATOPS MORNING REPORT – January 8, 2001

POES

No change in the status of the POES spacecraft.

NOAA-14 rev 31031 / F at 1342Z on January 6:  L5B noisy from spacecraft
and 96.3% recovered.  Another recovery attempt no possible due to
spacecraft recorder sequence management constraints.

NOAA-14 rev 31039 / F at 0302Z on January 7:  G5B noisy from spacecraft
and 86.4% of the expected data transferred to CEMSCS with 417 data gaps. 
Another recovery attempt no possible due to spacecraft recorder sequence
management constraints.

NOAA-14 rev 31042 / W at 0831Z on January 7:  L5B noisy from spacecraft
and 96.2% of the expected data transferred to CEMSCS with 27 data gaps. 
Another recovery attempt no possible due to spacecraft recorder sequence
management constraints.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

At 1857Z and 1942Z on January 5:  GOES-8 image frame breaks due
to high bit error rate.

At 1302Z and 1310Z on January 6:  GOES-8 image frame breaks due
to high bit error rate.

At 1949Z on January 6:  GOES-10 image frame break due to SPS ground
software reset.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.